# A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v1

@@ 3.1 General case
Figure. 1: The braid $B^{\#}\in B_{n+2}$. The box with ‘$B$’ in the center represents the braid $B$ on $n$ strands.
Associated to $B$, consider a braid $B^{\#}$ on $(n+2)$ strands defined by
$B^{\#}\coloneqq\sigma_{n}\cdot(\sigma_{n+1})^{3}\cdot B\in B_{n+2}.$
See Figure 1. The closure of $B^{\#}$ gives a knot which is the connected sum $K\#3_{1}$ of $K$ and the right-handed trefoil knot $3_{1}$.
In general, $HC^{\mathrm{ab}}_{0}(K\#3_{1})$ is a complicated $R$-algebra. However, it has the following property.
@@
We define an $R$-coefficient polynomial $F\in R[X]$ by
$F(X)\coloneqq X^{2}+(\mu+1-\mu^{-1}U)X+(\mu-1).$
Then, there exists a unital $R$-algebra map
$J\colon R[X]/(F)\to HC^{\mathrm{ab}}_{0}(K\#3_{1}).$
@@
We compute $\phi_{B^{\#}}(a_{n+2,*})$ by (10). Since $\phi_{\sigma_{k}}(a_{n+2,*})=a_{n+2,*}$ for every $k\in\{1,\dots,n-1\}$, $\phi_{B}(a_{n+2,*})=a_{n+2,*}$. Furthermore,
${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{B}}$${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+2,*}-a_{n+2,n+1}a_{n+1,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}+a_{n+1,n+2}(a_{n+2,*}-a_{n+2,n+1}a_{n+1,*})}$$\scriptstyle{\phi_{\sigma_{n}}}$${a_{n,*}+a_{n,n+2}(a_{n+2,*}-a_{n+2,n}a_{n,*}).}$
Let us write $P=a_{n,n+2}$ and $Q=a_{n+2,n}$. Then, the bottom $(n+2)$-th row of $\bm{\Phi}^{L}_{B^{\#}}$ is
$\begin{pmatrix}0&\cdots&0&1-PQ&0&P\end{pmatrix}.$
It follows that the $(n+2,n)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$ is
$\widehat{\mathbb{A}}_{n+2,n}-((1-PQ)\mathbb{A}_{n,n}+P\mathbb{A}_{n+2,n})$ $=-\mu Q-((1-PQ)(1-\mu)+P(-\mu Q))$
$=PQ-\mu Q-(1-\mu),$
@@
and the $(n+2,n+2)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$ is
$\widehat{\mathbb{A}}_{n+2,n+2}-((1-PQ)\mathbb{A}_{n,n+2}+P\mathbb{A}_{n+2,n+2})$ $=(U-\mu)-((1-PQ)P+P(1-\mu))$
$=PQP-(2-\mu)P+(U-\mu).$
Hence $P^{2}Q^{2}-(2-\mu)PQ+(U-\mu)Q\in\mathcal{I}_{B^{\#}}$ and by substituting $PQ$ with $\mu Q+(1-\mu)$, we can find an element of $\mathcal{I}_{B^{\#}}$
$(\mu Q+(1-\mu))^{2}-(2-\mu)(\mu Q+(1-\mu))+(U-\mu)Q,$
which is equal to
$\mu^{2}Q^{2}+(2\mu(1-\mu)-(2-\mu)\mu+(U-\mu))Q+(1-\mu)^{2}-(2-\mu)(1-\mu)$
=\ = $\mu^{2}Q^{2}+(-\mu^{2}-\mu+U)Q+(\mu-1).$
Therefore, if we define $F(X)=X^{2}+(\mu+1-\mu^{-1}U)X+(\mu-1)$, we have a well-defined unital $R$-algebra map
$J\colon R[X]/(F)\to HC^{\mathrm{ab}}_{0}(K\#3_{1})$
which sends $X$ to $-\mu Q=-\mu a_{n+2,n}$. ∎
Let us check this property in the simplest case.
@@
From a braid representation $\sigma_{1}^{3}\in B_{2}$ of $3_{1}$, it is not difficult to compute $HC_{0}(3_{1})$ as outlined in [16, Exercise 3.10]. $HC_{0}(3_{1})$ is isomorphic to a commutative $R$-algebra $R[Y]/(G_{1},G_{2})$, where
$\begin{array}[]{cc}G_{1}(Y)\coloneqq UY^{2}-\mu UY+\lambda\mu^{3}(1-\mu),&G_{2}(Y)\coloneqq UY^{2}+\lambda\mu^{2}Y+\lambda\mu^{2}(\mu-U).\end{array}$
See [16, Exercise 5.5]. We have identities
$YG_{1}+\mu G_{2}$ $=UY^{3}+\lambda\mu^{3}(2-\mu)Y+\lambda\mu^{3}(\mu-U),$
$UY(YG_{1}+\mu G_{2})$ $=(UY^{2})^{2}+\lambda\mu^{3}(2-\mu)(UY^{2})+\lambda\mu^{3}(\mu-U)UY,$
$UY^{2}$ $=G_{1}+(\mu UY-\lambda\mu^{3}(1-\mu)),$
so the ideal $(G_{1},G_{2})$ has an element
$(\mu UY-\lambda\mu^{3}(1-\mu))^{2}+\lambda\mu^{3}(2-\mu)(\mu UY-\lambda\mu^{3}(1-\mu))+\lambda\mu^{3}(\mu-U)UY$
=\ = $(\mu UY)^{2}+(\lambda\mu^{4}+\lambda\mu^{3}-\lambda\mu^{2}U)(\mu UY)+(\lambda\mu^{3})^{2}(\mu-1).$
This means that a unital $R$-algebra map
$R[X]/(F)\to HC^{\mathrm{ab}}_{0}(3_{1})\cong R[Y]/(G_{1},G_{2})\colon X\mapsto\lambda^{-1}\mu^{-2}UY=(\lambda\mu^{3})^{-1}(\mu UY)$
is well-defined for the polynomial $F$ in Lemma 3.6, since
$(\lambda\mu^{3})^{2}F(\lambda^{-1}\mu^{-2}UY)=(\mu UY)^{2}+(\lambda\mu^{3})(\mu+1-\mu^{-1}U)(\mu UY)+(\lambda\mu^{3})^{2}(\mu-1)\in(G_{1},G_{2}).$
@@
If we reduce the coefficient ring $R$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$, Lemma 3.6 implies the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map
$\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(\left.F\right|_{U=1})\to\left.HC^{\mathrm{ab}}_{0}(K\#3_{1})\right|_{U=1}$
induced by $J$. (Here, the notation $\left.HC_{0}(K\#3_{1})\right|_{U=1}$ for an $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra is the same one as in [16, Section 4].) However, this assertion is trivial. Indeed,
$\left.F\right|_{U=1}(X)=X^{2}+(\mu+1-\mu^{-1})X+(\mu-1)=(X+\mu)(X+1-\mu^{-1})$
and thus, for any unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra $A$, there exists a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(\left.F\right|_{U=1})\to A\colon X\mapsto-\mu$.
Lemma 3.6 implies the following constraint on augmentations.

### v2

@@ 3.2.2 On augmentations of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$ for $m\geq 0$
Figure. 1: The braid $B^{\#}_{m}\in B_{n+2}$ for $m=1$. The box with ‘$B$’ in the center represents the braid $B$ on $n$ strands.
By using (10), let us compute $\phi_{B^{\#}_{m}}(a_{n+2,*})=\left(\phi_{\sigma_{n}}\circ\phi_{(\sigma_{n+1})^{2m+1}}\circ\phi_{B}\right)(a_{n+2,*})$. First, since $\phi_{\sigma_{k}}(a_{n+2,*})=a_{n+2,*}$ for every $k\in\{1,\dots,n-1\}$, $\phi_{B}(a_{n+2,*})=a_{n+2,*}$. Next, we compute $\phi_{(\sigma_{n+1})^{2m+1}}(a_{n+2,*})$ by the induction on $m=0,1,\dots$.
@@
Let $\hat{f}_{m}$ and $\hat{g}_{m}$ be elements of the free ring $\mathbb{Z}\langle a_{n+1,n+2},a_{n+2,n+1}\rangle$ for every $m\in\mathbb{Z}_{\geq 0}$ determined inductively on $m=0,1,\dots$ by $\hat{f}_{0}=1$, $\hat{g}_{0}=0$, and
$\begin{cases}\hat{f}_{m+1}=\hat{f}_{m}\cdot(1-a_{n+1,n+2}a_{n+2,n+1})-\hat{g}_{m}\cdot a_{n+2,n+1},\\ \hat{g}_{m+1}=\hat{f}_{m}\cdot a_{n+1,n+2}+\hat{g}_{m}.\end{cases}$ (14)
Then, for every $m\in\mathbb{Z}_{\geq 0}$,
$\phi_{(\sigma_{n+1})^{2m+1}}(a_{n+2,*})=\hat{f}_{m}\cdot a_{n+1,*}+\hat{g}_{m}\cdot a_{n+2,*}.$ (15)
@@
We prove by induction on $m$. Since $\phi_{\sigma_{n+1}}(a_{n+2,*})=a_{n+1,*}$, (15) for $m=0$ is satisfied by setting $\hat{f}_{0}=1$ and $\hat{g}_{0}=0$. Suppose that (15) holds for some $m\in\mathbb{Z}_{\geq 0}$. For readability, let us rewrite
$\begin{array}[]{cccc}a_{n+1,n+2}=p,&a_{n+2,n+1}=q,&\hat{f}_{m}=\hat{f}_{m}(p,q),&\hat{g}_{m}=\hat{g}_{m}(p,q).\end{array}$
Note that $\phi_{\sigma_{n+1}}$ maps $p$ to q-q and $q$ to p-p. Let $\hat{f}_{m}(-q,-p),\hat{g}_{m}(-q,-p)\in\mathbb{Z}\langle p,q\rangle$ be the elements obtained from $\hat{f}_{m}(p,q)$ and $\hat{g}_{m}(p,q)$ by replacing $p$ with q-q and $q$ with p-p. Then, we can compute that
$\phi_{(\sigma_{n+1})^{2m+2}}(a_{n+2},*)$ $=\phi_{\sigma_{n+1}}\left(\hat{f}_{m}(p,q)\cdot a_{n+1,*}+\hat{g}_{m}(p,q)\cdot a_{n+2,*}\right)$
$=\hat{f}_{m}(-q,-p)\cdot(a_{n+2,*}-q\cdot a_{n+1,*})+\hat{g}_{m}(-q,-p)\cdot a_{n+1,*}$
$=\left(-\hat{f}_{m}(-q,-p)q+\hat{g}_{m}(-q,-p)\right)\cdot a_{n+1,*}+\hat{f}_{m}(-q,-p)\cdot a_{n+2,*},$
$\phi_{(\sigma_{n+1})^{2m+3}}(a_{n+2},*)$ $=\left(-\hat{f}_{m}(p,q)(-p)+\hat{g}_{m}(p,q)\right)\cdot(a_{n+2,*}-q\cdot a_{n+1,*})+\hat{f}_{m}(p,q)\cdot a_{n+1,*}$
$=\left(\hat{f}_{m}(p,q)(1-pq)-\hat{g}_{m}(p,q)q\right)\cdot a_{n+1,*}+\left(\hat{f}_{m}(p,q)p+\hat{g}_{m}(p,q)\right)\cdot a_{n+2,*}.$
Therefore, if we set $\hat{f}_{m+1}$ and $\hat{g}_{m+1}$ by (14), then (15) with $m$ replaced by $m+1$ holds. ∎
We apply $\phi_{\sigma_{n}}$ to (15), then
$\phi_{B^{\#}_{m}}(a_{n+2,*})$ $=\left(\phi_{\sigma_{n}}\circ\phi_{(\sigma_{n+1})^{2m+1}}\circ\phi_{B}\right)(a_{n+2,*})$
$=\left(\phi_{\sigma_{n}}\circ\phi_{(\sigma_{n+1})^{2m+1}}\right)(a_{n+2,*})$
$=\phi_{\sigma_{n}}(\hat{f}_{m})\cdot a_{n,*}+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot a_{n+2,*}.$
Since $\phi_{\sigma_{n}}$ maps $a_{n+1,n+2}$ to $a_{n,n+2}$ and $a_{n+2,n+1}$ to $a_{n+2,n}$, the elements $\phi_{\sigma_{n}}(\hat{f}_{m}),\phi_{\sigma_{n}}(\hat{g}_{m})\in\mathbb{Z}\langle a_{n,n+2},a_{n+2,n}\rangle$ are obtained from $\hat{f}_{m}$ and $\hat{g}_{m}$ by replacing $a_{n+1,n+2}$ with $a_{n,n+2}$ and $a_{n+2,n+1}$ with $a_{n+2,n}$.
It follows that the bottom ($n+2$)-th row of $\bm{\Phi}^{L}_{B^{\#}_{m}}$ is
$\begin{pmatrix}0&\cdots&0&\phi_{\sigma_{n}}(\hat{f}_{m})&0&\phi_{\sigma_{n}}(\hat{g}_{m})\end{pmatrix}$.
@@
Let $c^{m}_{i,j}\in\mathcal{A}_{n+2}$ denote the $(i,j)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}_{m}}\cdot\mathbb{A}$. We focus on the $(n+2,n)$-th and $(n+2,n+2)$-th entries:
$\begin{split}c^{m}_{n+2,n}&=\widehat{\mathbb{A}}_{n+2,n}-\left(\phi_{\sigma_{n}}(\hat{f}_{m})\cdot\mathbb{A}_{n,n}+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot\mathbb{A}_{n+2,n}\right)\\ &=-\mu a_{n+2,n}-\left(\phi_{\sigma_{n}}(\hat{f}_{m})\cdot(1-\mu)+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot(-\mu a_{n+2,n})\right),\\ c^{m}_{n+2,n+2}&=\widehat{\mathbb{A}}_{n+2,n+2}-\left(\phi_{\sigma_{n}}(\hat{f}_{m})\cdot\mathbb{A}_{n,n+2}+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot\mathbb{A}_{n+2,n+2}\right)\\ &=(U-\mu)-\left(\phi_{\sigma_{n}}(\hat{f}_{m})\cdot a_{n,n+2}+\phi_{\sigma_{n}}(\hat{g}_{m})\cdot(1-\mu)\right).\end{split}$ (16)
Using the map $\pi_{n+2}$ of (13), let us define elements of $R[a_{i,j}\mid 1\leq i,j\leq n+2,\ i\neq j]$ by
$\begin{array}[]{cc}f_{m}\coloneqq\pi_{n+2}(\phi_{\sigma_{n}}(\hat{f}_{m})),&g_{m}\coloneqq\pi_{n+2}(\phi_{\sigma_{n}}(\hat{f}_{m})),\end{array}$
and denote $X=a_{n,n+2}$ and $Y=a_{n+2,n}$. By (14), $f_{m}$ and $g_{m}$ are polynomials in $\mathbb{Z}[X,Y]$ determined inductively by $f_{0}=1$, $g_{0}=0$ and
$\begin{array}[]{cc}f_{m+1}=(1-XY)f_{m}-Yg_{m},&g_{m+1}=Xf_{m}+g_{m},\end{array}$ (18)
for every $m\in\mathbb{Z}_{\geq 0}$. Let us also define $F_{m},G_{m}\in\mathbb{Z}[\mu][X,Y]$ by
$\begin{array}[]{cc}F_{m}\coloneqq(1-\mu)f_{m}-\mu Yg_{m},&G_{m}\coloneqq Xf_{m}+(1-\mu)g_{m}.\end{array}$
Then, $\pi_{n+2}(c^{m}_{n+2,n})=-\mu Y-F_{m}$ and $\pi_{n+2}(c^{m}_{n+2,n+2})=U-\mu-G_{m}$ by (16).
Lemma 3.9.
Let us define $h_{m},k_{m}\in\mathbb{Z}[\mu][T]$ inductively on $m=0,1,\dots$ by $h_{0}=1-\mu$, $k_{0}=1$ and
$\begin{array}[]{cc}h_{m+1}=h_{m}-T\cdot k_{m},&k_{m+1}=h_{m}+(1-T)\cdot k_{m},\end{array}$ (20)
Then, for every $m\in\mathbb{Z}_{\geq 0}$,
$\begin{array}[]{cc}F_{m}(X,Y)=h_{m}(XY),&G_{m}(X,Y)=X\cdot k_{m}(XY).\end{array}$ (22)
Proof.
We prove by induction on $m$. Since $F_{0}=1-\mu$ and $G_{0}=X$, (22) for $m=0$ is satisfied by setting $h_{0}\coloneqq 1-\mu$ and $k_{0}\coloneqq 1$. Suppose that (22) holds for some $m\in\mathbb{Z}_{\geq 0}$. Using (18), we compute that
$F_{m+1}$ $=(1-\mu)f_{m+1}-\mu Yg_{m+1}$
$=(1-\mu)((1-XY)f_{m}-Yg_{m})-\mu Y(Xf_{m}+g_{m})$
$=(1-\mu-XY)f_{m}-Yg_{m}$
$=((1-\mu)f_{m}-\mu Yg_{m})-Y(Xf_{m}+(1-\mu)g_{m})$
$=F_{m}-YG_{m}$
$=h_{m}(XY)-XY\cdot k_{m}(XY),$
$G_{m+1}$ $=Xf_{m+1}+(1-\mu)g_{m+1}$
$=X((1-XY)f_{m}-Yg_{m})+(1-\mu)(Xf_{m}+g_{m})$
$=X(2-\mu-XY)f_{m}+(1-\mu-XY)g_{m}$
$=X((1-\mu)f_{m}-\mu Yg_{m})+(1-XY)(Xf_{m}+(1-\mu)g_{m})$
$=XF_{m}+(1-XY)G_{m}$
$=X\left(h_{m}(XY)+(1-XY)\cdot k_{m}(XY)\right).$
Therefore, if we set $h_{m+1},k_{m+1}\in\mathbb{Z}[\mu][T]$ by (20), then (22) with $m$ replaced by $m+1$ holds. ∎
For every $m\in\mathbb{Z}_{\geq 0}$, let us define a polynomial $P_{m}\in\mathbb{Z}[\mu,U][T]\subset R[T]$ by
$P_{m}\coloneqq(U-\mu)\cdot h_{m}+\mu T\cdot k_{m}.$ (23)
Proposition 3.10.
Let $\mathbb{k}$ be a field. For any augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)\in(\mathbb{k}^{*})^{2}$, the polynomial $\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
Proof.
By definition, the ideal $\mathcal{I}_{B^{\#}_{m}}$ contains $\pi_{n+2}(c^{m}_{n+2,n})=-\mu Y-F_{m}$ and $\pi_{n+2}(c^{m}_{n+2,n+2})=U-\mu-G_{m}$. Moreover, we have an identity
$P_{m}(XY)$ $=(U-\mu)\cdot h_{m}(XY)+\mu XY\cdot k_{m}(XY)$
$=(U-\mu)\left(h_{m}(XY)+\mu Y\right)-\mu Y\left(U-\mu-X\cdot k_{m}(XY)\right)$
$=(U-\mu)(\mu Y+F_{m}(X,Y))-\mu Y(U-\mu-G_{m}(X,Y)),$
which implies that $P_{m}(XY)\in\mathcal{I}_{B^{\#}_{m}}$. Therefore, there exists a well-defined unital $R$-algebra map
$J_{m}\colon R[T]/(P_{m})\to R[a_{i,j}\mid 1\leq i,j\leq n+2,\ i\neq j]/\mathcal{I}_{B^{\#}_{m}}=HC^{\mathrm{ab}}_{0}(K\#T_{(2,2m+1)})$
which maps $T$ to $XY=a_{n,n+2}a_{n+2,n}$. Now the proposition follows from Lemma 3.4. ∎
@@
Let us check the existence of a map as $J_{m}$ in a simple case. We have $h_{1}=1-\mu-T$, $k_{1}=(1-\mu)+(1-T)=2-\mu-T$ and
$P_{1}(T)$ $=(U-\mu)(1-\mu-T)+\mu T(2-\mu-T)$
$=-\mu T^{2}+(3\mu-\mu^{2}-U)T+(U-\mu)(1-\mu).$
From a braid representation $(\sigma_{1})^{3}\in B_{2}$ of $T_{(2,3)}$, it is not difficult to compute $HC_{0}(T_{(2,3)})$ as outlined in [15, Exercise 3.10]. $HC_{0}(T_{(2,3)})$ is isomorphic to a commutative $R$-algebra $R[X]/(G_{1},G_{2})$, where
$\begin{array}[]{cc}G_{1}(X)\coloneqq UX^{2}-\mu UX+\lambda\mu^{3}(1-\mu),&G_{2}(X)\coloneqq UX^{2}+\lambda\mu^{2}X+\lambda\mu^{2}(\mu-U).\end{array}$
See [15, Exercise 5.5]. We have identities
$XG_{1}+\mu G_{2}$ $=UX^{3}+\lambda\mu^{3}(2-\mu)X+\lambda\mu^{3}(\mu-U),$
$UX(XG_{1}+\mu G_{2})$ $=(UX^{2})^{2}+\lambda\mu^{3}(2-\mu)(UX^{2})+\lambda\mu^{3}(\mu-U)UX,$
$\lambda\mu^{2}(\mu-U)G_{1}$ $=\lambda\mu^{2}(\mu-U)(UX^{2})-\lambda\mu^{3}(\mu-U)UX+\lambda^{2}\mu^{5}(1-\mu)(\mu-U),$
and
$UX(XG_{1}+\mu G_{2})+\lambda\mu^{2}(\mu-U)G_{1}$
$=\$$(UX^{2})^{2}+\lambda\mu^{2}(3\mu-\mu^{2}-U)(UX^{2})+\lambda^{2}\mu^{5}(1-\mu)(\mu-U)$
[…]
Therefore, we have a well-defined unital $R$-algebra map
$R[T]/(P_{1})\to HC^{\mathrm{ab}}_{0}(T_{(2,3)})\cong R[X]/(G_{1},G_{2})\colon T\mapsto-\lambda^{-1}\mu^{-3}UX^{2}.$
@@
Consider the reduction of the coefficient ring from $R$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$. By the substitution $U=1$, we have $\left.P_{0}\right|_{U=1}=(1-\mu)^{2}+\mu T$, and if we define $h^{\prime}_{m},k^{\prime}_{m}\in\mathbb{Z}[\mu^{\pm}]$ by $h^{\prime}_{m}\coloneqq\left.h_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$ and $k^{\prime}_{m}\coloneqq\left.k_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$, then
$\left.P_{m+1}\right|_{U=1,T=-\mu^{-1}(1-\mu)^{2}}$ $=(1-\mu)\cdot h^{\prime}_{m+1}+\mu(-\mu^{-1}(1-\mu)^{2})\cdot k^{\prime}_{m+1}$
$=(1-\mu)(h^{\prime}_{m}+\mu^{-1}(1-\mu)^{2}k^{\prime}_{m})-(1-\mu)^{2}(h^{\prime}_{m}+(1+\mu^{-1}(1-\mu)^{2})k^{\prime}_{m})$
$=\mu\left((1-\mu)h^{\prime}_{m})-(1-\mu)^{2}k^{\prime}_{m}\right)$
$=\mu\left.P_{m}\right|_{U=1,T=-\mu^{-1}(1-\mu)^{2}}.$
Here, the second equality follows from (20). By induction on $m=0,1,\dots$, we can see that the polynomial $\left.P_{m}\right|_{U=1}\in\mathbb{Z}[\mu^{\pm}][T]$ has a root $T=-\mu^{-1}(1-\mu)^{2}$. Hence, Proposition 3.10 is trivial for if an augmentation $\varepsilon$ satisfies $\varepsilon(U)=1$.
3.2.3 Application of $\mathbb{Q}$-valued augmentations
Notation. For every $q\in\mathbb{C}(T)$, we define its degree $\deg q\coloneqq\max\{\deg f,\deg g\}$, where $f,g\in\mathbb{C}[T]$ are coprime polynomials such that $g\neq 0$ and $q=\frac{f}{g}$.
Remark 3.13.
Let us take $q=\frac{f}{g}\in\mathbb{C}(T)$ as above. Then, $\deg q\in\mathbb{Z}_{\geq 0}$ coincides with the degree of the continuous map $\hat{q}\colon\mathbb{C}P^{1}\to\mathbb{C}P^{1}$ such that $\hat{q}([z:1])=[f(z):g(z)]$ for every $z\in\mathbb{C}$ and
$\hat{q}([1:0])=\begin{cases}[\lim_{z\to\infty}\frac{f(z)}{g(z)}:1]&\text{ if }\deg f\leq\deg g,\\ [1:0]&\text{ if }\deg f>\deg g.\end{cases}$
Consider $r=\frac{f^{\prime}}{g^{\prime}}\in\mathbb{C}(T^{\prime})$ for $f^{\prime},g^{\prime}\in\mathbb{C}(T^{\prime})$ such that $\left.g^{\prime}\right|_{T^{\prime}=q}\neq 0$ in $\mathbb{C}(T)$. Then, the degree of $\left.r\right|_{T^{\prime}=q}\in\mathbb{C}(T)$ is computed by $\deg r\cdot\deg q$, which is the degree of the composite map $\hat{r}\circ\hat{q}\colon\mathbb{C}P^{1}\to\mathbb{C}P^{1}$.
In the following, we fix $y_{0}\in\mathbb{Q}^{*}\setminus\{1\}$ (one may choose, for instance, $y_{0}=2$). By the substitution $\mu=y_{0}$, let us define $\bar{h}_{m},\bar{k}_{m}\in\mathbb{Q}[T]$ as
$\begin{array}[]{cc}\bar{h}_{m}\coloneqq\left.h_{m}\right|_{\mu=y_{0}},&\bar{k}_{m}\coloneqq\left.k_{m}\right|_{\mu=y_{0}}.\end{array}$